### v2

@@ 4.6. RQ6: How effective are the DSLs?
A huge number of 29 DSLs illustrates the security DSL on smaller, typically to some extent realistic, yet still artificial examples. However, only for 17 DSLs these examples are actually part of an empirical evaluation. In these cases, the examples make up for a benchmark suite or are used in experiments with users of the DSL. In the other 12 DSLs, the examples are mainly a vehicle to illustratively explain the DSL to the reader but are sold as an empirical evaluation.
@@
The main vehicle for empirical evaluation is the case study (42 DSLs). Thereby, the main objective of a case study is to assess the feasibility of the DSL for the intended use case in its natural context (Runeson and Höst, 2009). To this end, the DSL is applied to one or multiple case studies and discussed in the paper. On average, the papers in this category contain 2.5 case studies, with a maximum of 12 case studies, resulting in a median of 1 case study. The case studies themselves are real-world systems or realistic systems of varying sizes. External experts are involved only rarely.
@@
Seven user studies were conducted to evaluate the concrete usage of the security DSL by target stakeholders or in the concrete application context. The studies typically aimed to assess the usability for stakeholders (4 DSLs), the applicability to real-world contexts (4 DSLs), and the understandability (2 DSLs) and usability (1 DSL) of the DSL.
@@
Four DSLs are evaluated by applying them to some sort of benchmark. Benchmarks are typically created for competitive evaluation in a specific domain and allow comparison of specific characteristics of the systems being evaluated (von Kistowski et al., 2015), in this case DSLs and their tools. However, only for two of them the entire benchmark suite is used, the other two selectively picked examples from benchmark suites.
@@
111122223333444455556666777788889999101010101111111112121212131313131414141415151515161616161717171718181818191919192020202021212121222222220044448888121212121616161620202020Number of DSLsexamplescase-studies
Figure 11. Number of investigated case-studies/examples
Since examples and case studies are the main methods used for evaluating security DSLs, in Fig. 11, we depict the number of case-studies and examples presented for the DSLs. The maximum number is 22 for one DSL, but on average only 3.2 examples or case studies per DSL are presented. The median is 2 DSLs. Altogether, the empirical evaluation of the security DSLs is mostly focused on individual cases that may be considered representative, but rigorous evaluation methods are often missing.
After identifying the method used for evaluation, we investigated what are the objectives of the evaluations. In our investigation of the evaluation of the security DSLs we identified 13 different objectives of the evaluation as well as the illustration of the DSL on examples which is presented as an evaluation of the DSL.
@@
The applicability to real-world systems and application scenarios has been evaluated for 9 DSLs in the corresponding publications. Thereby, the focus is primarily on validating that variations that might occur in practice are supported technically and not whether the proposed technique provides good results for all of these variations.
@@
Only 1 DSL is evaluated concerning the overhead it introduces for developers using the DSL. However, this is assessed by the authors in a case study they performed and not by an user study as one would expect.
@@
The effectiveness to actually achieve the intended security goal is evaluated for 8 DSLs. Effectiveness is evaluated via controlled experiments. On average, 9.3 more or less representative examples are used for this purpose, with the minimum being only one example and the maximum being 22 examples.
@@
Feasibility is usually evaluated on individual case studies by demonstrating that the proposed technique is feasible to solving the targeted problem. Corresponding to case studies being the mainly applied evaluation method, the majority of the security DSLs (40 DSLs) is evaluated concerning their principle feasibility for the intended task.
@@
The concrete performance of the security DSLs, typically focusing on the runtime of the corresponding tooling, is the evaluated for 18 DSLs.
@@
For one DSL the precision of a static analysis based on information expressed using the security DSL is evaluated.
@@
For 4 security DSLs the evaluation systematically focuses in experiments on how the DSLs scale for huge inputs.
@@
Whether stakeholders can easily understand what is expressed using the security DSLs is evaluated for 2 DSLs.
@@
The illustration of what the security DSL looks like to convey the idea of how it will be applied is based on examples for 12 security DSLs and is sold as a form of empirical evaluation. However, these examples are relatively small and may not fully align with the scope of a case study, which could potentially impact the validity of conclusions drawn from them.
Practical application
Only a minority of 14 security DSLs (12 %) has been applied in practice. Three of these DSLs are the OASIS standards SAML, XACML, and WS-SecurityPolicy that are widely used in practice. The mature academic security frameworks CORAS, STS, and UMLsec have been applied in various industrial context in projects and papers extending their original publications. All others have been applied to individual open-source projects or been evaluated together with industry partners.
Observation 6. Not only have most security DSLs rarely been applied in practice, their empirical validation is often lacking or of low quality, and even formal proofs are the exception. However, those DSLs that have been evaluated in practice or come from open source projects show a high interest from the industry.
@@ 5. Discussion
During our survey, we identified challenges related to the development and use of security DSLs as well as opportunities for future high-impact research. Overall, we developed a vision of the integrated utilization of security DSLs throughout the software development process as an effective means of developing secure systems. We first discuss identified challenges and opportunities, before outlining our vision and discussing why this integration will be an effective means.
@@ 5.1. Challenges & Opportunities
Although the use of security DSLs contributes to the security of a system, either by embedding security early on, i.e., during the requirements engineering or design phase, or by providing safe-to-use security measures during the implementation phase, the practical application seems to be limited. We identified few security DSLs that have been evaluated in practice, and even fewer DSLs that have become open standards. Investigating the barriers to practical adoption of security DSLs and the aspects that lead to successful adoption represents a significant research opportunity. The review of security DSLs and the findings from this survey are a first step in this direction.
Based on our extensive literature search, we can conclude that DSLs have already been presented for a wide range of security aspects and system components. We cannot identify a major security area where DSLs do not yet exist, but would obviously be helpful. While security DSLs exist for all major areas, there is still room for improvement in the details. For the access control domain, we observe that this detailed optimization is already well advanced, as DSLs have been explored for a variety of access control models. Also, for DSLs featuring information flow control, we see a trend toward more specific applications, such as concurrency or distributed systems.
@@
Most DSLs have been published in security community venues, but note that the security community itself is fragmented into several subdomains. The isolation of each DSL ultimately reflects this fragmentation. Unfortunately, we observe a low level of integration and combination. Most of the languages we examined produce only single artifacts and are not integrated into other frameworks. Nevertheless, there are some security DSLs that cover multiple phases of the SDLC and demonstrate the benefits of doing so, i.e., ensuring that planned security requirements are addressed as planned in later development phases. Further integration promises a massive increase in the security of systems under development.
The lack of editing support we observed is probably typical of scientific implementations, which focus on novelty rather than usability. Nevertheless, a few security DSLs have led to commercial products, enjoyed widespread adoption beyond the research community, or become an international standard. However, effective editing support is needed to put the DSLs into practice, and involves relevant research opportunities, particularly for the human-based security community, such as identifying what developers need to effectively use a security DSL.
@@
Particularly evident is, for scientific standards, the low level of evaluation for the DSLs we discovered. While the languages in the cryptography and information flow control domains usually provide strong formal security guarantees to undermine their effectiveness, most DSLs have not been evaluated with clear methodology. Many papers only demonstrate the feasibility of their language with small examples, while user studies have only been executed for a few DSLs. This is a call to arms to improve the evaluation of future work on security DSLs. However, this also provides an opportunity for empirical studies of the effectiveness of different security DSLs, including the potential to identify factors that influence the applicability or security gain of security DSLs.
@@ 5.2. Vision of an Integration Scenario
During our study, we identified various best practices and ideas that can help to effectively utilize security DSLs throughout the software development process. Based on these individual observations as well as the empirical data gathered, we developed a vision of deep integration of different security DSLs. To outline this vision and demonstrate how several of the DSLs examined in this work can be integrated to develop a secure system, we present a simple example of an EHR subsystem as motivated in the introduction. We map the DSLs to their respective phases in the SDLC, highlight how the DSLs help address specific security issues, and outline potential integrations. Our example consists of a measurement device, a computed tomography (CT) scanner, that sends its data over an open network to a server.
At the beginning of the development process, the (security) requirements are elicited and specified with the help of a security DSL, i.e., using STS and its supporting tools. The functional requirements of the system described above are modeled using a goal model notation oriented on SI* as shown in Fig. 6. Thereby, data planned to be exchanged between the individual entities is already explicitly specified on an abstract level. Here, the sensitivity of the measurement data gets identified and documented, for example, the CT images are labeled as security-critical data. Also using STS, basic authorization requirements for the intended usage scenarios are planned (see Fig. 12). For example, a doctor that is contacted as part of a remote consultation is allowed to read and modify health reports, while researchers are not allowed to have direct access. Thereby, checks provided as part of the editing support help in reasoning about the security requirements. For example, STS allows to propagate threats through the planned system to help identifying possible impacts. In Fig. 6, the tool shows by a red highlighting that the threat of the CT scanner being not reliable can propagate up to the treatment of a patient and might impact it.
@@
To prevent unintended leakage of information also at a lower level, Jif is used to implement parts of the data gathering logic. For example, a measurement data variable can be declared in Jif as shown in LABEL:lst:jif. The security label in the curly brackets would assure, that the measurement data that is generated by the scanner can only be read by the scanner itself and any entity that can act as AuthorizedUser. Furthermore, the measurement can only be modified by the scanner to ensure its integrity. This code written in Jif can be compiled to Java and thus seamlessly integrated with code written in Java. However, as for the design models, the data flow policies expressed in Jif must comply to the data flows planned during requirements engineering (cf. Figs. 6 and 12) as well as all further refinements, e.g, using security DSLs at design time. While we are not aware of any compliance check between data flow policies in Jif and ones expressed in other security DSLs, there are works that check whether design-time data flow contracts are properly implemented using static or dynamic analysis or by tailoring taint analyzers (Tuma et al., 2023; Peldszus et al., 2024; Lanzinger et al., 2024)
Listing 4: Excerpt from a data class with secure data flow constraints written in Jif.
@@
Finally, let us assume, the hospital participates in a medical study across multiple institutes, but privacy regulations prevent the disclosure of their patient’s measurement data. In response, secure multi party computation is applied, which allows computing functions, e.g., to determine average, median, and maximum values of the private data without revealing the data to other participating institutes. The involved parties can implement the functionality in SFDL 2.0 without cryptographic knowledge and the FairplayMP framework generates the desired code. As Jif, SFDL 2.0 compiles to Java bytecode that can be executed on the JVM, therefore allowing implementation-level integration.
Figure 12. Authorization view for an electronic health care system in STS.
The existence of several information-flow-sensitive DSLs, such as Jif or SFDL 2.0, points to the final challenge in the outlined integration, which is to decide on the appropriate DSLs to select for fulfilling the individual security requirements. Here, the overview provided in this paper helps to manually select suitable security DSLs for individual tasks. However, suitable security DSLs for the implementation of measures such as MPC should be automatically proposed as part of the development process, just as UMLsec was used to identify the required cryptography (cf. Fig. 8). Integrating such techniques into the reasoning of design-time security DSLs is the first step. The second step should be to automatically propose the most appropriate DSL that implements all the required features and best fits the design choices of the system.

### v3

@@ 4.6.1. Empirical Methods
29 DSLs are illustrated on smaller examples, which are often realistic, but still artificial. However, only for 17 these examples are actually part of an empirical evaluation. In these cases, the examples make up for a benchmark suite or are used in experiments with users of the DSL. In the other 12 DSLs, the examples are mainly a vehicle to illustratively explain the DSL to the reader but are sold as an empirical evaluation.
@@
The main vehicle for empirical evaluation is the case study (42 DSLs). Thereby, the main objective of a case study is to assess the feasibility of the DSL for the intended use case in its natural context (Runeson and Höst, 2009). To this end, the DSL is applied to one or more case studies and the application (and experiences) are reported in the paper. On average, the papers in this category contain 2.5 case studies, with a maximum of 12 case studies, resulting in a median of 1 case study. The case studies themselves are real-world systems or realistic systems of varying sizes. External experts are involved only rarely.
@@
Seven user studies were conducted to evaluate the concrete usage of the DSL by target stakeholders or in the concrete application context. The studies typically aim to assess the usability for stakeholders (4 DSLs), the applicability to real-world contexts (4 DSLs), and the understandability (2 DSLs) and usability (1 DSL).
@@
Four DSLs are evaluated based on benchmarks. Benchmarks are typically created for competitive evaluation in a specific domain and allow comparison of specific characteristics of the systems being evaluated (von Kistowski et al., 2015). However, only for two of the four DSLs the entire benchmark suite is used, the other two selectively picked examples from benchmark suites.
@@
112233445566778899101011111212131314141515161617171818191920202121222204488121216162020Number of DSLsexamplescase-studies
Figure 11. Number of investigated case studies/examples
Since examples and case studies are the main methods used for evaluating DSLs, in Fig. 11, we depict the number of case studies and examples presented for the DSLs. The maximum number is 22 for one DSL, but on average only 3.2 examples or case studies per DSL are presented. The median is two. Altogether, the empirical evaluation of the DSLs is mostly focused on individual cases that may be considered representative, but rigorous evaluation methods are often missing.
4.6.2. Formal Methods
The publications of 40 DSLs (33 %) state some form of security guarantee. We observed a strong correlation between the availability of an attacker model and a security guarantee: 79 % of the DSLs with an attacker model have a security guarantee and explicitly state the security objective. On the other hand, DSLs with a focus on access control do not have an attacker model (81 %) and do not provide a security guarantee in most cases (93 %). In general, when a security guarantee is provided by a DSL, it should be confirmed by some form of evaluation. Since a significant number of DSLs have a formal proof, we provide some details on this aspect.
Not applicable.:
67 % of our DSLs do not come with a formal proof, since formal proof techniques are not applicable to them, e.g., due to their informal nature. Only 33 % of the DSLs provide a security guarantee, for example, many languages in the field of information flow control assert a form of non-interference, a concept introduced by Goguen and Meseguer (Goguen and Meseguer, 1982), which implies that the observable output of a system must be independent of its sensitive input.
Formal proof.:
Out of the set of 30 DSLs with a security guarantee, 73 % provide a formal proof that assures the correctness of the claimed guarantee. These are mostly from domains with strong formal foundations, e.g., cryptography or information flow control.
No proof.:
Still, 12 DSLs remain that contain security aspects that could be accompanied by a formal proof, but no formal proof is provided. This are 10% of all DSLs and 40 % of the 30 DSL for which a formal proof would be applicable.
4.6.3. Objectives
After identifying the method used for evaluation, we investigated what are the objectives of the evaluations. We identified 13 different objectives of the evaluation as well as the illustration of the DSL on examples, which is presented as an evaluation of the DSL.
@@ 4.6.3. Objectives
Nine DSLs are evaluated concerning their applicability to real-world systems and application scenarios, i.e., whether the assumptions of the scientific worlds meet practice. Thereby, the focus is primarily on validating that variations that might occur in practice are supported technically and not whether the proposed technique provides good results for all variations.
@@
One DSL is evaluated concerning the overhead it introduces for developers using the DSL. However, this is assessed by the authors themselves in a case study they conducted, not by a user study as one might expect.
@@
The effectiveness to actually achieve the intended security goal is evaluated for 8 DSLs. Effectiveness is evaluated via controlled experiments. On average, 9.3 more or less representative examples are used for this purpose, with the minimum being only one example and the maximum being 22 examples. Note that the effectiveness of a DSL is not limited to concrete security guarantees; indirect security benefits are also relevant. For example, by simplifying the representation of an instantiation, a DSL helps to detect or avoid errors.
@@
Feasibility is usually evaluated on individual case studies by demonstrating that the proposed technique is feasible to solving the targeted problem. Thereby, the main focus is to demonstrate that the target problem is solvable using the DSL in some setting, which not necessarily has to be generalizable or entirely realistic. Corresponding to case studies being the mainly applied evaluation method, the majority of the DSLs (40 DSLs) is evaluated concerning their principle feasibility for the intended task.
@@
The concrete performance of the DSLs, typically focusing on the runtime of the corresponding tooling, is the evaluated for 18 DSLs.
@@
For one DSL the precision of a static analysis is evaluated.
@@
Four DSLs are evaluated how they scale for huge inputs.
@@
Whether stakeholders can easily understand what is expressed using the DSLs is evaluated for 2 DSLs.
@@
The illustration of what the DSL looks like to convey the idea of how it will be applied is based on examples for 12 DSLs and is sold as a form of empirical evaluation. However, these examples are relatively small and may not fully align with the scope of a case study, which could potentially impact the validity of conclusions drawn from them.
4.6.4. Practical Application
Only a minority of 14 DSLs (12 %) report on experiences from applications in practice. Three of these DSLs are the OASIS standards SAML, XACML, and WS-SecurityPolicy that are widely used in practice. The mature academic security frameworks CORAS, STS, and UMLsec have been applied in various industrial context in projects and papers extending their original publications. All others have been applied to individual open-source projects or been evaluated together with industry partners.
Observation 6. Not only have most security DSLs rarely been applied in practice, their empirical validation is often lacking or of low quality. However, in some domains, a significant share of DSLs offer formal security guarantees and those DSLs that have been evaluated in practice or come from open source projects show a high interest from the industry.
@@ 5. Discussion
We now discuss the challenges we identified related to the development and use of security DSLs, opportunities for future research, and a vision for integrating security DSLs throughout the SDLC.
5.1. Towards an Evaluation Framework
Despite analyzing a large number of security DSLs, we found few reports of their practical usage in the scientific literature. Instead, we observed an unsystematic empirical evaluation of security DSLs, underscoring the need for a more systematic assessment. To start the construction of a general evaluation framework for assessing security DSLs, Table 3 aggregates the details on the formal and empirical evaluation of security DSLs, as well as aspects identified in the general literature on DSL quality inspired by the Goal-Question-Metric paradigm for defining and interpreting software measurement (Basili, 1992). The goals were derived from the aspects addressed in RQ6 focusing on the evaluation of security DSLs (see Sec. 4.6), removing aspects that are claimed to be part of an evaluation but do not come with measurable metrics or qualitative arguments, i.e., Illustrate. For every goal, we extracted relevant questions to answer when assessing a security DSL as well as suitable metrics or qualitative arguments for answering the questions. When multiple aspects extracted from the papers containing the evaluation of a DSL were merged into one goal, the individual aspects are covered by the questions. Thereafter, we triangulated with general literature on usability aspects of DSLs and evaluating DSLs, extending goals, questions, and metrics with additional ones observed in the literature (annotated with references in Table 3). Particularly relevant was the qualitative DSL assessment framework (Kahraman and Bilgen, 2015) and the usability evaluation taxonomy (Rodrigues et al., 2017). We assessed every aspect presented in the literature with respect to the specifics of security DSLs, considering also insights from multiple perspectives, e.g., reusability of artifacts written in a DSL is considered in a framework as its own category (Kahraman and Bilgen, 2015), but a user study shows that it did not contribute to the success of the DSL (Hermans et al., 2009). Finally, we performed a sanitation step, merging closely related aspects into a single aspect, e.g., considering Scalability as a subaspect of Performance.
The investigated literature on security DSLs has focused on technical aspects such as correctness, effectiveness, and performance, while more general works on DSL qualities do not explicitly consider correctness and how it is ensured. Yet, these aspects are paramount in the domain of security. While the literature on security DSLs also discusses functional suitability and productivity, relevant aspects for practical use, such as usability, are barely discussed. Existing quality literature demonstrates the breadth of the spectrum that must be considered.
Table 3. Towards an evaluation framework: Goals, questions, and methods for assessing security DSLs
Goals Questions Methods Metrics/Qualitative Arguments
Compatibility*
(Kahraman and Bilgen, 2015; Hermans et al., 2009)
How compatible is the DSL with domain-specific (development) process? Case Studies, Examples SDLC phases(Kahraman and Bilgen, 2015)
[…]
Correctness Is the DSL’s implementation correct in respect to the security aspect expressed? Formal Proofs n/a
How understandable is the security DSL for stakeholders? User Studies # correctly understood concepts, user rating, semiotic clarity
Effectiveness (Barisic et al., 2011; Rodrigues et al., 2017; Hermans et al., 2009) How effective is the DSL in reaching its security goal? Benchmarks, Experiments recall (# vulnerabilites found/prevented)(Hermans et al., 2009), precision (# false positives)
Case Studies # vulnerabilites found/prevented
How effective is the DSL in facilitating security engineering tasks? Experiment, Case Studies # changes needed, # LoC needed/saved
How good is the quality reached by the security DSL? Benchmarks, Experiments size of output (i.e., LoC, # gates, chip area, power consumption)
Extensibility*
(Kahraman and Bilgen, 2015; Barisic et al., 2012)
[…]
Is the DSL appropriate to solve the intended problem? Examples, Case Studies problem solved
Is the security DSL applicable to real-world systems and scenarios? Case Studies practical constrains met, barriers to adoption
Integrability*
(Kahraman and Bilgen, 2015)
[…]
Performance How much computation resources are needed by the tooling of the DSL? Benchmarks, Experiments execution time(Mohagheghi and Haugen, 2010), memory usage, storage usage
[…]
How does the runtime performance scale? Benchmarks, Experiments scalability for large/complex cases
[…]
Usability
(Kahraman and Bilgen, 2015; Barisic et al., 2011, 2012) (Rodrigues et al., 2017; Hermans et al., 2009)
[…]
* Derived from additional literature and not observed in the papers/material on the security DSLs
@@ 5.2. Challenges & Opportunities
Although security DSLs clearly contribute to the security of a system, the low number of documented practical evaluations indicate low rate of adoption. Investigating the barriers to practical adoption of security DSLs represents a significant research opportunity.
Barriers to adoption that we have identified are primarily a lack of availability and usability. For two-thirds of the DSLs, the framework presented in the publications is not publicly available. The prevalent primitive level of usability, e.g., due to missing editing support for the DSL, is certainly a limitation for practical adoption. We assume that the focus on novelty and theoretical foundations in academia in combination with a lack of incentives for practicality and maintenance are the cause for this situation. Besides this, particularly design-time security DSLs require specific artifacts, e.g., models, that are often not available in practice (Gorschek et al., 2014). However, industrial threat modeling frameworks such as Microsoft’s STRIDE significantly rely on design models, i.e, data flow diagrams (Shostack, 2008). Nevertheless, a few security DSLs have led to commercial products, enjoyed widespread adoption beyond the research community, or become an international standard. However, effective editing support is needed to put the DSLs into practice, and involves relevant research opportunities, particularly for the human-based security community, such as identifying what developers need to effectively use a security DSL. In line with this, an initial study focusing on UMLsec (Ebad, 2022) shows that developers consider the prototypical state of the tooling, i.e., leading to a limited user experience, and the required training in using the DSL to be main reasons for the lack of adoption.
An objective, standardized evaluation framework would help to compare security DSLs, e.g., with respect to their strength and weaknesses, and assist practitioners in identifying the most appropriate one for their needs. Our initial framework for assessing security DSLs (cf. Table 3) and our supplementary material (Krausz et al., 2025), consisting of detailed labeling per DSL, could be used as a basis for developing this framework. As the categories of security DSLs identified in this work address different domains and problems, a framework tailored to each category may be required. In this regard, as related work (see Sec. 7), we discuss existing surveys that address individual categories.
Our literature review shows that DSLs exist for nearly all major security aspects and system components, leaving few gaps. Access control DSLs, such as SAML, WS-SecurityPolicy, and XACML, are widely used and considered mature, with limited need for further extensions. DSLs for information flow control and secure computation increasingly target specialized contexts, such as concurrency, distributed systems, or MPC. However, despite the importance of testing systems for security, we did not observe any DSLs specific to writing security tests or the secure deployment of systems. While there are many general DSLs for specifying test cases, such as Robot (Bisht, 2013), and even sub-DSLs within it (Peldszus et al., 2023), security-focused DSLs appear to be absent. Perhaps existing DSLs are sufficient for specifying security tests and require further automation to generate them from other DSLs, for example. Conversely, current studies show that security testing is still primarily manual, (Hermann et al., 2025a), e.g., based on pen testing, which might indicate a lack of need for security-specific testing DSLs.
@@
Although most DSLs have been published in security venues, the security community itself is fragmented into several subdomains. The isolation of each DSL ultimately reflects this fragmentation. Unfortunately, we observe a low level of integration and combination. Most of DSLs produce only single artifacts and are not integrated into other frameworks. Nevertheless, there are some security DSLs that cover multiple phases of the SDLC and demonstrate the benefits of doing so, i.e., ensuring that planned security requirements are addressed as planned in later development phases.
@@
Particularly evident is, for scientific standards, the low level of evaluation for the DSLs we discovered. While the languages in the cryptography and information flow control domains usually provide strong formal security guarantees to undermine their effectiveness, most DSLs have not been evaluated with clear methodology. Many papers only demonstrate the feasibility of their language with small examples, while user studies have only been executed for a few DSLs. This provides an opportunity for empirical studies of the effectiveness of different security DSLs and to identify factors that influence the applicability or security gain of security DSLs.
In practice, one of the main challenges in developing secure systems is the planning of appropriate security features and their secure realization (Hermann et al., 2025b). While some categories of security DSLs usually come with tool support for generating secure code or compiling instances into executable code, these are primarily DSLs targeting custom cryptographic security features or enriching code with information flow control policies. However, both still require developers to manually embed the security features into a system. Although security DSLs have been used for code generation (Arogundade et al., 2020; Peldszus, 2022), these approaches mainly generate structural code around security features rather than securely embedding the features themselves. Therefore, it only prevents a security feature from being missing altogether. Emerging technologies, such as generative AI, have been used to automate code generation, but have security issues (Pearce et al., 2022). Using DSLs as input to generative AI could be an effective building block for generating secure code, including the secure use of security features.
@@ 5.3. Vision of an Integration Scenario
During our study, we identified various best practices and ideas that can help to effectively utilize security DSLs throughout software development. To demonstrate how several of the DSLs examined in this work can be integrated to develop a secure system, we designed a simple example of an EHR subsystem as motivated in the Introduction. Specifically, we mapped the DSLs to their respective phases in the SDLC, and we then highlighted how the DSLs help to address specific security issues in order to outline potential integrations. This example consists of a measurement device (a computed tomography (CT) scanner) that sends its data over an open network to a server.
At the beginning of the development process, (security) requirements are elicited and specified with STS and its supporting tools. The functional requirements of the system described above are modeled using a goal model notation oriented on SI* (see Fig. 6). Thereby, data planned to be exchanged is already explicitly specified on an abstract level. Here, the sensitivity of the measurement data gets identified and documented, for example, the CT images are labeled as security-critical data. Also using STS, basic authorization requirements for the intended usage scenarios are planned (see Fig. 12). Thereby, checks provided as part of editing support help in reasoning about the security requirements. For example, STS allows to propagate threats through the planned system to help identifying possible impacts. In Fig. 6, the tool highlights in red that the threat of the CT scanner being not reliable can propagate the treatment of a patient and might impact it.
@@
To prevent unintended information leakage at a lower level, Jif is used to implement data gathering logic. For example, a measurement data variable can be declared in Jif as shown in LABEL:lst:jif. The security label in the curly brackets assures, that data generated by the scanner can only be read by the scanner itself and any entity that can act as AuthorizedUser. Furthermore, the measurement can only be modified by the scanner to ensure its integrity. This code can be compiled to Java and thus seamlessly integrated with code written in Java. However, as for the design models, the data flow policies expressed in Jif must comply to the data flows planned during requirements engineering (cf. Figs. 6 and 12) as well as all further refinements, e.g, using security DSLs at design time. While we are not aware of any compliance check between data flow policies in Jif and ones expressed in other security DSLs, there are works that check whether design-time data flow contracts are properly implemented using static or dynamic analysis or by tailoring taint analyzers (Tuma et al., 2023; Peldszus et al., 2024; Lanzinger et al., 2024)
Listing 4: Excerpt from a data class with secure data flow constraints written in Jif
@@
To allow participating in medical studies across multiple institutes, while avoiding disclosure of patient measurements due to privacy regulations, secure multi party computation is used to handle statistical requests, e,g., determining median values of patient measurements. This can be implemented in SFDL 2.0 with little cryptographic knowledge and FairplayMP generates code, therefore allowing implementation-level integration. The existence of several information-flow-sensitive DSLs, such as Jif or SFDL 2.0, points to the final challenge in the outlined integration, which is to decide on the appropriate DSLs to select for fulfilling the individual security requirements. Here, the overview provided in this paper helps to manually select suitable security DSLs for individual tasks.
Figure 12. Authorization view for an electronic health care system in STS
Note that we selected the DSLs so that integration and combination are already feasible or would only require reasonable effort. In general, combining multiple DSLs, can lead to significant complexity due to semantic (Neto et al., 2013) and technical (Horst and Rumpe, 2013) differences. Integrability and compatibility are also featured as quality aspects of our proposed evaluation framework. For integrating independent components, different types of reuse need to be considered, such as type, implementation, and binding hierarchies (Mezini and Ostermann, 2002). Composing multiple DSLs requires creating interfaces for composing term representations and defining interpretations across languages (Hofer and Ostermann, 2010). The necessary steps for systematically integrating DSLs and their common concepts are open research questions (Heithoff et al., 2023; Pfeiffer et al., 2023).